OpenStudy (anonymous):

standard form : containing ( 2, -5) and ( 4, 4 )

OpenStudy (anonymous):

i assume you mean a line in y=mx+b containing (2, -5) and (4, 4)

OpenStudy (anonymous):

Ax+By=C

OpenStudy (anonymous):

Find the slope. Use Point slope with your information. Solve for y. Get the x term on to the same side as the y term by adding or subtracting it.

OpenStudy (aroub):

first find the slope 4+5/4-2 9/2 y=mx+b , m=9/2 , (4,4) 4=9/2(4)+b 4=36/4+b -36/4+16/4=b -5=b y=9/2x-5

OpenStudy (anonymous):

this would be applying rise over run, then plugging in a value, so m=(4+5)/(4-2)= 9/2 y=9/2*x+b 4=9/2*4+b 4=18+b b=-14 y=9/2*x-14

OpenStudy (anonymous):

aroub, yours cannot be right, x=2 gets you 4, not -5

OpenStudy (anonymous):

you essed up when multiplying by 4, 9/2*4=36/2, not 36/4

OpenStudy (anonymous):

what does the little star mean?

OpenStudy (anonymous):

mulitply

OpenStudy (anonymous):

its common usage for multiplication, since x is used as a variable so often

OpenStudy (aroub):

oh yes ur right!! soorryy!! :(

OpenStudy (anonymous):

thanks yall(:

OpenStudy (anonymous):

9x-2y=28